BPJ stock is expected to earn 14.8 percent in a recession, 6.3 percent in a normal economy, and lose 4.7 percent in a booming economy. The probability of a boom is 20 percent while the probability of a normal economy is 55 percent. What is the expected rate of return on this stock?

A. 6.23%

B. 6.72%

C. 6.81%

D. 7.60%

E. 8.11%
Answer #1

Probability of recession=(100-20-55)=25%

expected rate of return on this stock=Respective returns*Respective probabilities

=(14.8*0.25)+(6.3*0.55)-(4.7*0.2)

=6.23%(Approx)